OpenStudy (anonymous):

find the exact solution log (x)+log (2x-1)=log 3 there all log little 3

OpenStudy (anonymous):

The base is same accros them so we can multiply the values: So: x(2x-1) = 3 2x^2-x =3 2x^2 - x - 3 = 0 Now calculate the x.

OpenStudy (anonymous):

(2x-3) (x+1)

OpenStudy (anonymous):

answers would be 3/2 and -1

OpenStudy (anonymous):

So x will be -1 and 3/2. remember that -1 is not acceptable because log of a negative number is undefined. So the only acceptable answer is 3/2

OpenStudy (anonymous):

thanks again cortez

OpenStudy (anonymous):

you're welcome, dont forget to mark as an answer if it helped you ;)